# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ 1.2.2 XVAE: A VAE incorporating the proposed max-id spatial model
The data reconstruction process relies on compactly supported local basis functions at pre-determined knot points, which are not updated with $\boldsymbol{\phi}_{d}$ of the decoder. Although one could choose the knots using a certain space-filling design, we propose a data-driven way to determine the number of knots, their locations, and the radius of basis functions as described in Section D.2 of the Supplementary Material, and show by simulation that this compares favorably to the XVAE initialized with the true knots/radii. Our XVAE implementation in R is publicly accessible on GitHub at https://github.com/likun-stat/XVAE.
1.3 Marginal and dependence properties
One major advantage of our approach over typical machine learning generative models is that we can explicitly derive the marginal and dependence properties of our spatial extremes model integrated within the VAE. In this section, we temporarily omit the subscript $t$ and examine the marginal and joint tail behavior of the model (3) at one specific time point. For notational simplicity, we write $X_{j}=X(\boldsymbol{s}_{j})$, $\omega_{kj}=\omega_{k}(\boldsymbol{s}_{j})$, $k=1,\ldots,K$, $j=1,\ldots,n_{s}$, and define $\mathcal{C}_{j}=\{k:\omega_{kj}\neq 0,k=1,\ldots,K\}$. We require that any location $\boldsymbol{s}\in\mathcal{S}$ be covered by at least one basis function, thereby $\mathcal{C}_{j}$ cannot be empty for any $j$.
In spatial extremes, extremal dependence is commonly described by the measure
$\chi_{ij}(u)=\Pr\{F_{j}(X_{j})>u\mid F_{i}(X_{i})>u\}=\frac{\Pr\{F_{j}(X_{j})>u,F_{i}(X_{i})>u\}}{\Pr\{F_{i}(X_{i})>u\}}\in[0,1],$ (9)
in which $u\in(0,1)$ is a (high) threshold and $F_{i}$ and $F_{j}$ are the continuous marginal distribution functions for $X_{i}$ and $X_{j}$, respectively. When $u$ is close to one, $\chi_{ij}(u)$ quantifies the probability that one variable is extreme given that the other variable is similarly extreme. If $\chi_{ij}=\lim_{u\rightarrow 1}\chi_{ij}(u)=0$, $X_{i}$ and $X_{j}$ are said to be asymptotically independent (AI), and if $\chi_{ij}=\lim_{u\rightarrow 1}\chi_{ij}(u)>0$, $X_{i}$ and $X_{j}$ are asymptotically dependent (AD).
1.3.1 Marginal distributions
To examine $\chi_{ij}=\lim_{u\rightarrow 1}\chi_{ij}(u)$, we first study the marginal distributions of the process (3).
Proposition 1.1.
Let $\mathcal{D}=\{k:\gamma_{k}=0,\;k=1,\ldots,K\}$ and $\bar{\mathcal{D}}$ be the complement of $\mathcal{D}$. For process (3),the marginal distribution function of $X_{j}=X(\boldsymbol{s}_{j})$ can be written as
$F_{j}(x)=\exp\left\{\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left(\gamma_{k}+{\tau}^{\frac{1}{\alpha_{0}}}\omega_{kj}^{\frac{1}{\alpha}}x^{-\frac{1}{\alpha_{0}}}\right)^{\alpha}\right\}.$ (10)
As $x\rightarrow\infty$, the survival function $\bar{F}_{j}(x)=1-F_{j}(x)$ satisfies
$\bar{F}_{j}(x)=c^{\prime}_{j}x^{-\frac{\alpha}{\alpha_{0}}}+c_{j}x^{-\frac{1}{\alpha_{0}}}+\left(d_{j}-\frac{c_{j}^{2}}{2}\right)x^{-\frac{2}{\alpha_{0}}}-\frac{c^{\prime}_{j}{}^{2}}{2}x^{-\frac{2\alpha}{\alpha_{0}}}-c^{\prime}_{j}c_{j}x^{-\frac{\alpha+1}{\alpha_{0}}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),$ (11)
where $c_{j}=\alpha\tau^{1/\alpha_{0}}\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha-1}\omega_{kj}^{1/\alpha}$, $c^{\prime}_{j}=\tau^{\alpha/\alpha_{0}}\sum_{k\in\mathcal{D}}\omega_{kj}$, and $d_{j}=\frac{\alpha(\alpha-1)}{2}\tau^{2/\alpha_{0}}\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha-2}\omega_{kj}^{2/\alpha}$.
The proof of this result can be found in Section B.2 of the Supplementary Material. It indicates that the process (3) has Pareto-like marginal tails at any location in the domain $\mathcal{S}$. If $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, that is, if the $j$th location is impacted by an “un-tilted knot” (i.e., a knot with $\gamma_{k}=0$ in the $\mathrm{expPS}(\alpha,\gamma_{k})$ distribution of the corresponding latent variable $Z_{k}$), then $\bar{F}_{j}(x)\sim c_{j}^{\prime}x^{-\frac{\alpha}{\alpha_{0}}}$ as $x\rightarrow\infty$ since $\alpha\in(0,1)$. If, however, the location is not within the reach of an un-tilted knot, then instead $\bar{F}_{j}(x)\sim c_{j}x^{-\frac{1}{\alpha_{0}}}$ as $x\rightarrow\infty$, which is less heavy-tailed.
The following result directly delineates how the quantile level changes as $u\rightarrow 1$.
Corollary 1.1.1.
As $t\rightarrow\infty$, the marginal quantile function $q_{j}(t)=F_{j}^{-1}(1-1/t)$ can be approximated as follows under the assumptions of Proposition 1.1:
$q_{j}(t)=\begin{cases}c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+\frac{\alpha_{0}c_{j}t^{1-1/\alpha}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}-\frac{\alpha_{0}t^{-1}}{2\alpha}+O\left(t^{-1/\alpha}\right)\right\},&\text{if }\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset,\\ c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+\alpha_{0}(\frac{d_{j}}{c_{j}^{2}}-\frac{1}{2})t^{-1}+o(t^{-1})\},&\text{if }\mathcal{C}_{j}\cap\mathcal{D}=\emptyset.\end{cases}$
The proof of this result can also be found in Section B.2 of the Supplementary Material. It will be used to derive the tail dependence structure for two arbitrary spatial locations.
1.3.2 Joint distribution
To derive the extremal dependence structure, we first calculate the joint distribution function of a $n_{s}$-variate random vector $(X_{1},\ldots,X_{n_{s}})^{T}$ drawn from the process (3).
Proposition 1.2.
Under the definitions and notation as established in the previous sections, for locations $\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n_{s}}\in\mathcal{S}$, the exact form of the joint distribution function of the random vector $(X_{1},\ldots,X_{n_{s}})^{T}$ can be written as
$F(x_{1},\ldots,x_{n_{s}})=\exp\left\{\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left(\gamma_{k}+\tau^{\frac{1}{\alpha_{0}}}\sum_{j=1}^{n_{s}}\omega_{kj}^{\frac{1}{\alpha}}x_{j}^{-\frac{1}{\alpha_{0}}}\right)^{\alpha}\right\}.$ (12)
The proof of Proposition 1.2 is given in Section B.3 of the Supplementary Material. Eq. (12) ensures that $F^{1/r}(x_{1},\ldots,x_{n_{s}})$ is a valid distribution function on $\mathbb{R}^{n_{s}}$ for any real $r>0$, of the same form as (12) but with tilting indices $\{\gamma_{1}/r^{1/\alpha},\ldots,\gamma_{K}/r^{1/\alpha}\}$ and scale parameter $\tau/r^{\alpha_{0}/\alpha}$. By definition, the process $\{X_{t}(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\}$ is thus max-infinitely divisible. It becomes max-stable only when it remains within the same location-scale family, i.e., when $\gamma_{1}=\cdots=\gamma_{K}=0$.
1.3.3 Tail dependence properties
We now characterize the tail dependence structure of process (3) using both $\chi_{ij}$ defined in Eq. (9) and the complementary measure $\eta_{ij}$ defined by $\Pr\{X_{i}>F_{i}^{-1}(u),X_{j}>F_{j}^{-1}(u)\}=\mathcal{L}\{(1-u)^{-1}\}(1-u)^{1/\eta_{ij}}$, where $\mathcal{L}$ is slowly varying at infinity, i.e., $\mathcal{L}(tx)/\mathcal{L}(t)\rightarrow 1$ as $t\rightarrow\infty$ for all $x>0$. The value of $\eta_{ij}\in(0,1]$ is used to differentiate between the different levels of dependence exhibited by an AI pair $(X_{i},X_{j})^{T}$. When $\eta_{ij}=1$ and $\mathcal{L}(t)\not\rightarrow 0$ as $t\rightarrow\infty$, $(X_{i},X_{j})^{T}$ is AD ($\chi_{ij}>0$), and the remaining cases are all AI ($\chi_{ij}=0$; see Ledford and Tawn, 1996), with stronger tail dependence for larger values of $\eta_{ij}$.
@@ 1.3.3 Tail dependence properties
Under the assumptions of Propositions 1.1 and 1.2, the process $\{X(\boldsymbol{s})\}$ defined in (3) has a tail dependence structure characterized as follows:
  1. (a)
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, we have $\chi_{ij}=0$ and $\eta_{ij}=1/2$.
  2. (b)
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have
    $t\Pr\{X_{i}>q_{i}(t),X_{j}>q_{j}(t)\}=\left\{\begin{matrix}c_{ij}t^{-\frac{1}{\alpha}}/(c_{i}c^{\prime}_{j}{}^{\frac{1}{\alpha}})+o(t^{-\frac{1}{\alpha}}),&\text{if }\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset,\\ t^{-1},&\text{if }\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset,\end{matrix}\right.$
    where $c_{ij}=\alpha(\alpha-1)\tau^{2/\alpha_{0}}\sum_{k\in\mathcal{C}_{i}\cap\mathcal{C}_{j}}\theta_{k}^{\alpha-2}\omega_{ki}^{1/\alpha}\omega_{kj}^{1/\alpha}$. This leads to $\chi_{ij}=0$, $\eta_{ij}=\frac{\alpha}{\alpha+1}$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset$ and $1/2$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$.
  3. (c)
    If $\mathcal{C}_{i}\cap\mathcal{D}\neq\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have
    $t\Pr\{X_{i}>q_{i}(t),X_{j}>q_{j}(t)\}=2-d_{ij}-({c_{i}}/{c^{\prime}_{i}{}^{1/\alpha}}+{c_{j}}/{c^{\prime}_{j}{}^{1/\alpha}})t^{1-\frac{1}{\alpha}}-O(t^{2-\frac{2}{\alpha}}),$
    where $d_{ij}=\tau^{{\alpha}/{\alpha_{0}}}\sum_{k\in\mathcal{D}}\{(\omega_{ki}/c^{\prime}_{i})^{1/\alpha}+(\omega_{kj}/c^{\prime}_{j})^{1/\alpha}\}^{\alpha}\in(1,2)$. Thus, $(X_{i},X_{j})^{T}$ is AD with $\chi_{ij}=2-d_{ij}$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset$, and AI with $\eta_{ij}=\alpha$ and $\chi_{ij}=0$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$.
The proof of this result can be found in Section B.4 of the Supplementary Material. The local dependence strength is proportional to the tail-heaviness of the latent variable of the closest knot. There is local AD if $\gamma_{k}=0$ and there is local AI if $\gamma_{k}>0$, as expected. Similar to Eq. (10), the sets $\mathcal{C}_{j}\cap\mathcal{D}$, $j=1,\ldots,n_{s}$, are crucial to the behavior of the so-called exponent function which occurs in the limiting distribution for normalized maxima (Huser and Wadsworth, 2019); see Remark 5 in Section B.4 for more discussion.
The compactness of the basis functions’ support yields long-range exact independence (thus, also AI) for two far-apart stations that are impacted by disjoint sets of basis functions; this is similar in spirit to the Cauchy convolution process of Krupskii and Huser (2022), though their model construction is different and less computationally tractable than ours.
2 Validation framework for extremes emulation
We propose a new validation framework that is tailored to assess skill in fitting both the full range of data and joint tail behavior in model outputs. This comprehensive framework can be used to evaluate the quality of emulations from any generative spatial extremes model and represents one of the contributions of this paper to the extremes literature.
2.1 Full range evaluation
To examine the quality of the emulation from the XVAE, we will predict at $n_{h}$ locations $\{\boldsymbol{h}_{i}:i=1,\ldots,n_{h}\}$ held out from the analyses. To perform these predictions, we calculate the basis function values at these locations, with which we can mix the encoded variables from Eq. (5) to get predicted values. For each time $t$ and holdout location $\boldsymbol{h}_{i}$, denote the true observation of $X_{t}(\boldsymbol{h}_{i})$ by $x_{it}$ and the emulated prediction by $x^{*}_{it}$. Then the mean squared prediction error (MSPE) for time $t$ is
$\mathrm{MSPE}_{t}=\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}(x_{it}-x^{*}_{it})^{2},$
where $t=1,\ldots,n_{t}$. All MSPEs from different time replicates can be summarized in a boxplot; see Section 3 for example. Similarly, we can calculate the continuously ranked probability score (CRPS; Matheson and Winkler, 1976; Gneiting and Raftery, 2007) across time for each location, i.e.,
where $F_{i}$ is the marginal distribution estimated using parameters at the holdout location $\boldsymbol{h}_{i}$, $i=1,\ldots,n_{h}$, and again $x^{*}_{it}$ is the emulated value. Smaller CRPS indicates that the distribution $F_{i}$ is concentrated around $x^{*}_{it}$, and thus can be used to measure how well the distribution fits all emulated values. Section 3 also shows how we present the CRPS values from all holdout locations for each emulation. In addition, we will examine the quantile-quantile (QQ)-plots obtained by pooling the spatial data into the same plot to check if the spatial input and the emulation have similar ranges and quantiles.
2.2 Empirical tail dependence measures
To assess the tail dependence structure of the emulated fields, we will estimate $\chi_{ij}(u)$ defined in Eq. (9) empirically in two ways. First, to examine the overall dependence strength, we treat $\{X(\boldsymbol{s})\}$ as if it had a stationary and isotropic dependence structure so that $\chi_{ij}(u)\equiv\chi_{h}(u)$, with $h=||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||$ being the distance between locations. Then for a fixed $h$, we find all pairs of locations with similar distances (within a small tolerance, say $\epsilon=0.001$), and compute the empirical conditional probabilities $\widehat{\chi}_{h}(u)$ at a grid of $u$ values. Confidence envelopes can be calculated by regarding the outcome (i.e., simultaneously exceed $u$ or not) of each pair as a Bernoulli variable and computing pointwise binomial confidence intervals, assuming that all pairs of points are independent from each other. Examples in Section 3 demonstrate how this empirical measure can be used to compare the extremal dependence structures between the spatial data input and realizations from the emulator. While this metric does not completely characterize the non-stationarity in the process, it is still well-defined as a summary statistic and carries important information about the average decay of dependence with distance irrespective of the direction.
Second, to avoid the stationary assumption, we can choose a reference point denoted by $\boldsymbol{s}_{0}$ and estimate the pairwise $\chi_{0j}(u)$ empirically between $\boldsymbol{s}_{0}$ and all observed locations $\boldsymbol{s}_{j}$ in the spatial domain $\mathcal{S}$. These pairwise estimates can then be presented using a raster plot (if gridded) or a heat plot. Section 4 shows examples of the empirical $\chi_{0j}(u)$, $u=0.85$, estimated from the real and emulated datasets, where $\boldsymbol{s}_{0}$ is the center of $\mathcal{S}$.
2.3 Areal radius of exceedance
We further propose a tail dependence coefficient that formally summarizes the overall dependence strength over the entire spatial domain. This metric characterizes the spatial extent of extreme events conditional on an arbitrary reference point in the domain (e.g., the center of $\mathcal{S}$) exceeding a particular quantile $u$. Zhang et al. (2023) formulated the metric on an empirical basis and named it the averaged radius of exceedances (ARE).
Given a large number of independent replicates (say $n_{r}$) from $\{X(\boldsymbol{s})\}$ on a dense regular grid $\mathcal{G}=\{\boldsymbol{g}_{i}\in\mathcal{S}:i=1,\ldots,n_{g}\}$ over the domain $\mathcal{S}$ with side length $\psi>0$, denote the replicates by $\boldsymbol{X}_{r}=\{X_{r}(\boldsymbol{g}_{i}):i=1,\ldots,n_{g}\}$, $r=1,\ldots,n_{r}$. The empirical marginal distribution functions at $\boldsymbol{g}_{i}$ can then be obtained as $\hat{F}_{i}(x)=n_{r}^{-1}\sum_{r=1}^{n_{r}}\mathbbm{1}(X_{ir}\leq x)$, where $X_{ir}=X_{r}(\boldsymbol{g}_{i})$ and $\mathbbm{1}\{\cdot\}$ is the indicator function. We then transform $(X_{i1},\cdots,X_{in_{r}})^{T}$ to the uniform scale via $U_{ir}=\hat{F}_{i}(X_{ir})$, $r=1,\ldots,n_{r}$. Let $\boldsymbol{U}_{r}=\{U_{ir}:i=1,\ldots,n_{g}\}$ and $U_{0r}=\hat{F}_{0}\{X_{r}(\boldsymbol{s}_{0})\}$. The ARE metric at the threshold $u$ is defined by
The summation $\psi^{2}\sum_{i=1}^{n_{g}}\mathbbm{1}(U_{ir}>u,U_{0r}>u)$ in Eq. (13) calculates the area of all grid cells exceeding the extremeness level $u$ at the same replicate $r$ as the reference location $\boldsymbol{s}_{0}$; dividing it by $\pi$ and taking the square root thus yields the “radius” of a circular exceedance region that has the same spatial extent. Additionally, Eq. (13) averages over all replicates with the reference location exceeding the extremeness level $u$. Therefore, $\widehat{\mathrm{ARE}}_{\psi}(u)$ has the same units as $\psi$, or the distance metric used on the domain $\mathcal{S}$, which makes it a more straightforward metric for domain scientists because it reflects the average length scale of the extreme events (e.g., warm pool size in SST data).
If there is an infinite number of independent replicates, the following result ensures that $\widehat{\mathrm{ARE}}_{\psi}(u)$ almost surely converges to
$\mathrm{ARE}_{\psi}(u)=\left(\psi^{2}\sum_{i=1}^{n_{g}}\chi_{0i}(u)/\pi\right)^{1/2},$ (14)
where $\chi_{0i}(u)$ is here the $\chi$-measure between locations $\boldsymbol{s}_{0}$ and $\boldsymbol{g}_{i}$ defined in Eq. (9).
Theorem 2.1.
@@ 2.3 Areal radius of exceedance
For a fixed regular grid $\mathcal{G}$ with side length $\psi$, a reference location $\boldsymbol{s}_{0}$ and $u\in(0,1)$, we have $\widehat{\mathrm{ARE}}_{\psi}(u)\rightarrow\mathrm{ARE}_{\psi}(u)$ almost surely as $n_{r}\rightarrow\infty$.
@@
From this result, we see that $\widehat{\mathrm{ARE}}_{\psi}(u)$ and its limit, which do not require stationarity or isotropy, quantify the square root of spatial average of $\chi_{0i}(u)$. Due to the presence of the white noise term $\epsilon(\boldsymbol{s})$ form, there is no version of the process $\{X(\boldsymbol{s})\}$ that has measurable paths, which means that $X(\boldsymbol{s})\not\rightarrow X(\boldsymbol{s}_{0})$ (in probability) as $\boldsymbol{s}\rightarrow\boldsymbol{s}_{0}$. However, from Theorem 1.3, we know that there is continuity in the dependence measure $\chi_{0i}$ because $\epsilon(\boldsymbol{s})$ has little impact on the dependence structure of the mixture $Y(\boldsymbol{s})$. That is, $\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}$, denoting the $\chi$-measure between location $\boldsymbol{s}_{0}$ and $\boldsymbol{s}\in\mathcal{S}$, is a continuous function of $\boldsymbol{s}\in\mathcal{S}$ when fixing the reference location $\boldsymbol{s}_{0}$; we define this property to be tail-continuous for $\boldsymbol{s}_{0}$. The following result further confirms that under the tail-continuity, $\widehat{\mathrm{ARE}}_{\psi}(u)$ also converges to the square root of spatial integration of $\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}$ as $u\rightarrow 1$ and as $\mathcal{G}$ becomes infinitely dense.
@@
Tail-continuity is met by many spatial extremes models, like max-stable, inverted-max-stable, and others (e.g., Opitz, 2016; Huser and Wadsworth, 2019; Wadsworth and Tawn, 2022). Our model (3) also adheres to tail-continuity, as indicated by Theorem 1.3.
@@
Together, Theorems 2.1 and 2.2 ensure that $\widehat{\mathrm{ARE}}_{\psi}(u)\approx\left\{\int_{\mathcal{S}}\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}\mathrm{d}\boldsymbol{s}\right\}^{1/2}/\pi^{1/2}$ if there are a large number of replicates from the process $\{X(\boldsymbol{s})\}$ on a very dense grid $\mathcal{G}$.
Similarly, we can estimate $\mathrm{ARE}_{\psi}(u)$ for the emulator by running the decoder repeatedly to obtain emulated replicates of $\{X(\boldsymbol{s})\}$ on the same grid. By comparing the $\mathrm{ARE}_{\psi}(u)$ estimates at a series of $u$ levels, we can evaluate whether the spatially-aggregated exceedance is consistent between the spatial data inputs and their XVAE emulation counterparts.
2.4 Uncertainty quantification
The decoder (8) functions as a neural estimator for $(\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}$. Examining its inferential power is crucial, as accurate emulation heavily relies on precise characterization of spatial inputs. Drawing a substantial number of samples from the variational distribution $q_{\boldsymbol{\phi}_{e}}(\cdot\mid\boldsymbol{x}_{t})$ (which is close to ${p_{\boldsymbol{\phi}_{d}}(\cdot\mid\boldsymbol{x}_{t})}$; recall Section 1.1) allows us to obtain Monte Carlo estimates of the dependence parameters $(\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}$ using the decoder (8). Aggregating these estimates provides approximate samples of the posterior $(\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}\mid\{\boldsymbol{x}_{t}:t=1,\ldots,n_{t}\}$, which enables the calculation of point estimates (posterior mean or maximum a posteriori) and the construction of approximate confidence regions for uncertainty quantification.

### v4

@@ 3.2 XVAE: A VAE incorporating the proposed max-id model
The data reconstruction process relies on compactly supported local basis functions at pre-determined knot points, which are not updated with $\boldsymbol{\phi}_{d}$ of the decoder. Although one could choose the knots using a certain space-filling design, we propose a data-driven way to determine the number of knots, their locations, and the radius of basis functions as described in Section D.4 of the Supplementary Material. We show by simulation that this compares favorably to the XVAE initialized with the true knots/radii. Our XVAE implementation in R is publicly accessible on GitHub at https://github.com/likun-stat/XVAE.
Uncertainty quantification:
The decoder (11) functions as a neural estimator for $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}$. Examining its inferential power is crucial, as accurate emulation heavily relies on precise characterization of spatial inputs. Drawing a substantial number of samples from the variational distribution $q_{\boldsymbol{\phi}_{e}}(\cdot\mid\boldsymbol{x}_{t})$ (which is close to ${p_{\boldsymbol{\phi}_{d}}(\cdot\mid\boldsymbol{x}_{t})}$; recall Section 2.2) allows us to obtain Monte Carlo estimates of the dependence parameters $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}$ using the decoder (11). Combining these estimates yields an approximate sample from the posterior, $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}\mid\{\boldsymbol{x}_{t}:t=1,\ldots,n_{t}\}$, which enables the calculation of point estimates (posterior mean or maximum a posteriori) and approximate confidence regions for UQ.
3.3 Validation framework for extremes emulation
We propose a validation framework tailored to assess both the full data range and the joint tail behavior in outputs from any generative spatial extremes model.
First, we predict at held-out locations and calculate the mean squared prediction error (MSPE) and the continuous ranked probability score (CRPS; Matheson and Winkler, 1976; Gneiting and Raftery, 2007). Second, we estimate $\chi_{ij}(u)$, as defined in Eq. (1), using two methods: (1) To summarize the average decay of dependence with distance even if the process is non-stationary, we treat $\{X(\boldsymbol{s})\}$ as having a stationary, isotropic dependence structure, where $\chi_{ij}(u)\equiv\chi_{h}(u)$, with $h=||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||$ as the distance between locations. For a fixed $h$, we compute empirical conditional probabilities $\widehat{\chi}_{h}(u)$ across a grid of $u$ values; (2) To avoid the restrictive stationary working assumption, we select a reference point $\boldsymbol{s}_{0}$ and estimate the pairwise $\chi_{0j}(u)$ between $\boldsymbol{s}_{0}$ and other locations $\boldsymbol{s}_{j}$ in the spatial domain $\mathcal{S}$, which can be visualized using raster or heat plots. Third, we examine QQ-plots by pooling spatial data to compare the ranges and quantiles of the input and emulated field. Further details of these diagnostics are provided in Section B of the Supplementary Material.
Additionally, we propose using a novel joint tail dependence coefficient that formally summarizes the overall dependence strength over the entire spatial domain. This metric characterizes the spatial extent of extreme events conditional on an arbitrary reference point in the domain exceeding a particular quantile $u$. Zhang et al. (2023) formulated the metric on an empirical basis and named it the averaged radius of exceedances (ARE).
Given a large number of independent replicates (say $n_{r}$) from $\{X(\boldsymbol{s})\}$ on a dense regular grid $\mathcal{G}=\{\boldsymbol{g}_{i}\in\mathcal{S}:i=1,\ldots,n_{g}\}$ over the domain $\mathcal{S}$ with side length $\psi>0$, denote the replicates by $\boldsymbol{X}_{r}=\{X_{r}(\boldsymbol{g}_{i}):i=1,\ldots,n_{g}\}$, $r=1,\ldots,n_{r}$. The empirical marginal distribution functions at $\boldsymbol{g}_{i}$ can then be obtained as $\hat{F}_{i}(x)=n_{r}^{-1}\sum_{r=1}^{n_{r}}\mathbbm{1}(X_{ir}\leq x)$, where $X_{ir}=X_{r}(\boldsymbol{g}_{i})$ and $\mathbbm{1}\{\cdot\}$ is the indicator function. We then transform $(X_{i1},\ldots,X_{in_{r}})^{\rm T}$ to the uniform scale via $U_{ir}=\hat{F}_{i}(X_{ir})$, $r=1,\ldots,n_{r}$. Let $\boldsymbol{U}_{r}=\{U_{ir}:i=1,\ldots,n_{g}\}$ and $U_{0r}=\hat{F}_{0}\{X_{r}(\boldsymbol{s}_{0})\}$. The ARE metric at the threshold $u$ is defined by
The summation $\psi^{2}\sum_{i=1}^{n_{g}}\mathbbm{1}(U_{ir}>u,U_{0r}>u)$ in Eq. (12) calculates the area of all grid cells exceeding the extremeness level $u$ jointly with the reference location $\boldsymbol{s}_{0}$, for the same replicate $r$; dividing it by $\pi$ and taking the square root thus yields the “radius” of a circular exceedance region that has the same spatial extent. Additionally, Eq. (12) averages over all replicates with the reference location exceeding the extremeness level $u$. Therefore, $\widehat{\mathrm{ARE}}_{\psi}(u)$ has the same units as $\psi$, or the distance metric used on the domain $\mathcal{S}$, which makes it an interpretable metric for domain scientists because it reflects the average length scale of extreme events (e.g., warm pool size in SST data).
The following result ensures that $\widehat{\mathrm{ARE}}_{\psi}(u)$, which does not require stationarity or isotropy, converges to the square root of the spatial average of $\chi_{0i}(u)$ as $n_{r}\rightarrow\infty$.
@@ 3.3 Validation framework for extremes emulation
For a fixed regular grid $\mathcal{G}$ with side length $\psi$, a reference location $\boldsymbol{s}_{0}$ and $u\in(0,1)$, we have that, almost surely,
$\widehat{\mathrm{ARE}}_{\psi}(u)\rightarrow\mathrm{ARE}_{\psi}(u)=\left(\psi^{2}\sum_{i=1}^{n_{g}}\chi_{0i}(u)/\pi\right)^{1/2},$ (13)
as $n_{r}\rightarrow\infty$, where $\chi_{0i}(u)$ is the $\chi$-measure between locations $\boldsymbol{s}_{0}$ and $\boldsymbol{g}_{i}$ defined in Eq. (1).
@@
Due to the presence of the white noise $\{\epsilon(\boldsymbol{s})\}$, there is no version of the process $\{X(\boldsymbol{s})\}$ that has measurable paths, which means that $X(\boldsymbol{s})\not\rightarrow X(\boldsymbol{s}_{0})$ (in probability) as $\boldsymbol{s}\rightarrow\boldsymbol{s}_{0}$. Nevertheless, we know from Theorem 3.3 that there is continuity in the dependence measure $\chi_{0i}$ because $\{\epsilon(\boldsymbol{s})\}$ barely impacts the dependence structure of $\{Y(\boldsymbol{s})\}$. That is, $\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}$, denoting the $\chi$-measure between location $\boldsymbol{s}_{0}$ and $\boldsymbol{s}$, is a continuous function of $\boldsymbol{s}\in\mathcal{S}$ when fixing the reference location $\boldsymbol{s}_{0}$; we define this property as tail-continuity. The following result further confirms that under the tail-continuity, $\widehat{\mathrm{ARE}}_{\psi}(u)$ also converges to the square root of the spatial integral of $\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}$ as $u\rightarrow 1$ and as $\mathcal{G}$ becomes infinitely dense.
@@
Tail-continuity is met by many spatial extremes models, like max-stable, inverted-max-stable, and others (e.g., Opitz, 2016; Huser and Wadsworth, 2019; Krupskii and Huser, 2022). Our model (3) also adheres to tail-continuity, as indicated by Theorem 3.3.
@@
Together, Theorems 3.4 and 3.5 ensure that $\widehat{\mathrm{ARE}}_{\psi}(u)\approx\left\{\int_{\mathcal{S}}\chi_{\boldsymbol{s}_{0},\boldsymbol{s}}\mathrm{d}\boldsymbol{s}\right\}^{1/2}/\pi^{1/2}$ if there are a large number of replicates from the process $\{X(\boldsymbol{s})\}$ on a very dense grid $\mathcal{G}$.
Similarly, we can estimate $\mathrm{ARE}_{\psi}(u)$ for the emulator by running the decoder repeatedly to obtain emulated replicates of $\{X(\boldsymbol{s})\}$ on the same grid. By comparing the $\mathrm{ARE}_{\psi}(u)$ estimates at a series of $u$ levels, we can evaluate whether spatially-aggregated exceedances are consistent between the spatial data inputs and their XVAE emulation counterparts.